### v1

@@ 4 Signal denoising
To refine the normalization factors as proposed at the end of the previous section (cf. 21) we took 1000 Gaussian random signals with mean 00 and variance $1$. These signals were decomposed and the standard deviation for each of the detail components and for the approximation component was measured. For each of the detail components and the approximation component the average of these standard deviations was taken as the normalisation factor for the corresponding component. The magnitude of these normalization factors turned out to be around $1$.
To determine how the parameter $\theta$ influences the reconstruction results, we applied the Bayes method to the six wavelet test functions of the wavelet toolbox of Matlab [^1] We used Matlab version R2024B. Matlab is a registered trademark of the MathWorks.: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp and Mishmash, plotted in Fig. 1.
For each of these test functions we executed the Bayes method with soft thresholding for parameter values $\theta$ ranging from $0.1$ to $0.9$. The length of the input signals was $2^{17}$. For each value of $\theta$ $10$ samples of the input signal were considered and the average was taken.
Figure 3: Comparison of our denoising method with the ones of “wdenoise” of Matlab for the six wavelet test functions: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp, Mishmash.
In Fig. 2, the result is plotted for the six test functions. This figure shows the signal-to-noise ratio (SNR) of the input signal with respect to the SNR of the output signal. We recall that the SNR of the two signals $s_{1}$ and $s_{2}$ is defined as
@@
From the plots it is clear that the value $\theta=0.1$ is best or near-best for the 6 test functions. Hence, in the sequel, we fixed the value of $\theta$ to be equal to $0.1$.
In the next experiment, we compared our denoising method with the one of Matlab “wdenoise” (of the wavelet toolbox) for all choices of the wavelets that are available using this denoising function: bior1.1, bior1.3, bior1.5, bior2.2, bior2.4, bior2.6, coif1, coif2, coif3, coif4, coif5, db2, db3, db4, db5, db6, db7, db8, db9, db10, db11, rbio1.3, rbio1.5, rbio2.2, rbio2.4, rbio2.6, rbio2.8, sym2, sym3, sym4, sym5, sym6, sym7. Applying “wdenoise”, we used the default denoising method which is “Bayes” and soft thresholding. Also here we took the average of $10$ samples. Figure 3 shows the results for the 6 wavelet test functions.
@@ 5 Image compression
Image compression is a key application of wavelet transforms, aimed at reducing storage and transmission costs while maintaining the perceptual quality of the data (see, e.g., [3, 4, 1, 2, 16, 9], among many others). In this section, we illustrate the performance of VP wavelets for image compression, comparing them with traditional separable wavelets. It is important to clarify that we use here the term ”compression” somewhat loosely. Our approach involves only thresholding, without applying the full compression pipeline, which would typically include quantization and encoding. The goal of our study is in fact to assess the sparsifying ability - or compressibility - of the basis rather than to implement a complete compression scheme.
@@
We now present a comparative analysis of the reconstruction quality between the VP basis and the biorthogonal spline wavelet filters bior3.5. We selected such filters because they are widely used in image compression, thanks to their symmetry, compact support, and high number of vanishing moments. The evaluation is performed in terms of PSNR and SSIM, considering their variation as a function of $\log_{2}(\text{CR})$. For this comparison, we have fixed seven levels of compression ratio, namely $2^{k}$ for $k=1,\dots,7$. For a fair comparison, since the images have all a size of $512\times 512$, we have considered 4 levels of decomposition for VP and 7 levels for the bior3.5 filters, so that, at the very last level, a similar amount of approximation coefficients are retained (4 and 6, respectively). For the VP basis, a near-optimal $\theta$ parameter, providing nearly the best PSNR, was chosen from the interval $[0.5,0.8]$ based on the experimental tests illustrated above.

### v2

@@ 5 Signal denoising
Using the factors $\sqrt{3}$ and $\sqrt{3/2}$ as proposed at the end of the previous section already improved the results but we wanted to refine these normalization factors. Therefore, we took 1000 Gaussian random signals with mean 0 and variance $1$. These signals were decomposed and the standard deviation for each of the detail components and for the approximation component was measured. In case of orthogonal wavelets, the averages of these standard deviations would converge to one if the number of samples increases. For the non-orthogonal wavelets used in this paper, the magnitude of these averages turned out to be around $1$. We used these averages as normalization factors.
To determine how the parameter $\theta$ influences the reconstruction results, we applied the Bayes method to the six test signals of the wavelet toolbox of Matlab [^1] We used Matlab version R2024B. Matlab is a registered trademark of the MathWorks.: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp and Mishmash, plotted in Fig. 2.
For each of these test signals we executed the Bayes method with soft thresholding for parameter values $\theta$ ranging from $0.1$ to $0.9$. The length of the input signals was $2^{17}$. The depth of the iterations was $8$. For each value of $\theta$ $10$ samples of the input signal were considered and the average was taken as well as the standard deviaton with respect to the average.
Figure 4: Comparison of our denoising method with the ones of “wdenoise” of Matlab for the six test signals: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp, Mishmash.
In Fig. 3, the result is plotted for the six test signals. This figure shows the signal-to-noise ratio (SNR) of the input signal with respect to the SNR of the output signal. We recall that the SNR of the two signals $s_{1}$ and $s_{2}$ is defined as
@@
From the plots it is clear that the value $\theta=0.1$ is best or near-best for the 6 test signals. Hence, in the sequel, we fixed the value of $\theta$ to be equal to $0.1$. In Figure 5 the standard deviation divided by the average is plotted for $\theta=0.1$ in function of the input SNR (dB) and each of the test signals.
Figure 5: The standard deviation divided by the average in function of input SNR (dB) for each of the test function for $\theta=0.1$
In the next experiment, we compared our denoising method with the one of Matlab “wdenoise” (of the wavelet toolbox) for all choices of the wavelets that are available using this denoising function: bior1.1, bior1.3, bior1.5, bior2.2, bior2.4, bior2.6, coif1, coif2, coif3, coif4, coif5, db2, db3, db4, db5, db6, db7, db8, db9, db10, db11, rbio1.3, rbio1.5, rbio2.2, rbio2.4, rbio2.6, rbio2.8, sym2, sym3, sym4, sym5, sym6, sym7. Applying “wdenoise”, we used the default denoising method which is “Bayes” and soft thresholding. Also here we took the average of $10$ samples. Figure 4 shows the results for the 6 test signals.
@@ 6 Image compression
Image compression is a key application of wavelet transforms, aimed at reducing storage and transmission costs while maintaining the perceptual quality of the data (see, e.g., [1, 2, 3, 4, 6, 23, 13], among many others). In this section, we illustrate the performance of VP wavelets for image compression, comparing them with traditional separable wavelets. It is important to clarify that we use here the term ”compression” somewhat loosely. Our approach involves only thresholding, without applying the full compression pipeline, which would typically include quantization and encoding. The goal of our study is in fact to assess the sparsifying ability - or compressibility - of the basis rather than to implement a complete compression scheme.
@@
We now present a comparative analysis of the reconstruction quality between the VP basis and the biorthogonal spline wavelet filters bior3.5. We selected such filters because they are widely used in image compression, thanks to their symmetry, compact support, and high number of vanishing moments. The evaluation is performed in terms of PSNR and SSIM, considering their variation as a function of $\log_{2}(\text{CR})$. For this comparison, we have fixed seven levels of compression ratio, namely $2^{k}$ for $k=1,\dots,7$. For a fair comparison, since the images have all a size of $512\times 512$, we have considered 4 levels of decomposition for VP and 7 levels for the bior3.5 filters, so that, at the very last level, a similar amount of approximation coefficients are retained (4 and 6 coefficients, respectively). For the VP basis, a near-optimal $\theta$ parameter, providing nearly the best PSNR, was chosen from the interval $[0.5,0.8]$ based on the experimental tests illustrated above.
@@
Figure 9: Original boat image (top), full SSIM maps (middle), and reconstructed images (bottom) for the bior3.5 wavelet filter (left column) and the VP wavelet (right column).
Figure 10: Zoomed views of the SSIM maps (top row) and the reconstructed images (bottom row) for the bior3.5 wavelet filter (left column) and the VP wavelet (right column).
To illustrate the reconstruction quality visually, we show in Figs. 9 and 10 the results obtained for the ”Boats” image after compression at $CR=2^{4}$, retaining only about $6.25\%$ of the coefficients. The accompanying SSIM maps indicate the local structural similarity with the original image. Overall, the VP wavelets maintain sharper details. The zoomed views (Fig. 10) focus on one of the ropes within the boat, highlighting that VP reconstructions exhibit better preservation of fine structures. This visual evidence is consistent with the plot in Fig. 11, which shows, in logarithmic scale, the first 30000 wavelet coefficients (absolute values) sorted in decreasing order. The VP wavelet coefficients are smaller in magnitude than those of the classical transform, indicating that the VP decomposition captures most of the energy in the approximation subband and leaves only small residuals in the detail subband.
Figure 11: Absolute values (in log scale) of the first 30000 sorted wavelet detail coefficients for the biorthogonal (bor3.5) and VP decomposition applied to the Boats image
@@ Acknowledgments
The authors thank the anonymous referees for their helpful comments and suggestions, which improved the quality of the paper.